A local news agency would like to test the claim of the mayor that more than half the voters in the city support the construction. If p is the proportion of voters in favor of the construction, the news agency will test H 0 : p = .5 against H 1 : p < .5 to a .05 level of significance by surveying 500 voters. If 225 voters are in favor of the construction give the P-value and your conclusion.

Answer choices:

P=.0258, At least half the voters support construction.

P=.0258, Less than half the voters support construction.

P=.0129, Less than half the voters support construction.

P=.0129, At least half the voters support construction.
Answer #1

Solution:

Hypothesis are

H 0 : p = .5 against H 1 : p < .5

n = 500

x = 225

Let  \hat p be the sample proportion.

\hat p = x/n = 225/500 = 0.45

The test statistic z is

z =   Р-р р(1-р) п =  (0.45 - 0.5)/\sqrt{}[0.5*(1 - 0.5)/500] = -2.23

Now , observe that there is < sign in H1

So , test is LEFT TAILED.

For left tailed test ,

p value = P[Z < z] = P[Z < -2.23] = 0.0129

Given , significance level \alpha = 0.05

p value is less than \alpha

Reject H0

We can conclude that there is sufficient evidence to support the claim that  Less than half the voters support construction.

So, correct option is

P=.0129, Less than half the voters support construction.